The heights of a certain species of plant are normally distributed, with mean 22cm and standard deviation 3cm. What is the probability that a plant chosen at random will be between 19.75 and 24.25 cm tall?
I have no Idea on this.. Please help. thanks.

Answer by stanbon(75887)

The heights of a certain species of plant are normally distributed, with mean 22cm and standard deviation 3cm.
What is the probability that a plant chosen at random will be between 19.75 and 24.25 cm tall?
-------------------------------------------------
Find the z-scores of 19.75 and 24.25
z(19.75) = (19.75-22)/3 = -2.25/3 = -0.75
---
z(24.25) = (24.25-22)/3 = 2.25/3 = 0.75
---
P(19.75 < x < 24.25) = P(-.75 < z < 0.75) = 0.1295
===========================================================
Cheers,
Stan H.